OpenStudy (anonymous):

Divide please: 2x^3-x^2y-11xy^2+10y^3 / x-2y

OpenStudy (compassionate):

\[2x^3 - x^2y-11xy^2+10y^3 / x-2y\]

OpenStudy (anonymous):

that is correct

OpenStudy (compassionate):

Add like terms, can you do that?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

trying to figure out how to do this

OpenStudy (compassionate):

\[2^3 + 10x^3\] = ? \[-x^2y-11xy^2\]= ?

OpenStudy (anonymous):

the first one is 12x^3?

OpenStudy (anonymous):

-12x1y1 for the second one?

OpenStudy (compassionate):

No, 2^3 = 8, 8 + 10x^3 = ?

OpenStudy (anonymous):

18x^3

OpenStudy (compassionate):

Eh, I'm trying to get the point across that you can't add any of these :)

OpenStudy (anonymous):

oh

OpenStudy (compassionate):

\[\frac{ 2x^3-x^2y-11xy^2+10y^3 }{ x-2y }\] 2x - x - 5.5y + 5y^2

OpenStudy (anonymous):

so that would be 1x-5.5y+5y^2

OpenStudy (anonymous):

?

OpenStudy (compassionate):

Technically x-2y cannot be solved further, so the original expression is also the final expression.

OpenStudy (anonymous):

the one you wrote right? and no remainer? so that's my answer

OpenStudy (compassionate):

Well, yes and no. I'm still not sure what form of division you're doing. I know three different methods and applications. Which form of division is this?

OpenStudy (anonymous):

the problem just says to divide

OpenStudy (compassionate):

That's not very good advice... It doesn't specify what type of division.

OpenStudy (anonymous):

nope it just says Divide.

OpenStudy (anonymous):

and then for the answer box it has quotient =

OpenStudy (anonymous):

you want either synthetic division or long division. have you done either of these before?

OpenStudy (anonymous):

It has to be long divison because it dosen't say synthetic

OpenStudy (anonymous):

will u look up to the answer on up and tell me if that's right?

OpenStudy (anonymous):

ill have a go.. i have to admit i haven't tried algebra division with 2 variables

OpenStudy (anonymous):

i did it :D i think it factorises: